OpenStudy (anonymous):

Can someone solve...10^4 - 160 by facotring polynomials?

OpenStudy (anonymous):

Factoring*

OpenStudy (anonymous):

10^4 = 2^4*5^4 and 160 = 2^5*5 10^4 - 160 = 2^4*5^4 - 2^5*5...

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

its right tho?